# Beyond Trend and Periodicity: Guiding Time Series Forecasting with Textual Cues

### v2

@@ I.2 Dataset Structure
We get the weather time series data from: https://www.bgc-jena.mpg.de/wetter/ and weather report from https://www.timeanddate.com/weather/germany/jena/historic. We will publish this dataset with CC BY-NC-SA 4.0 licence since the data source forbids commercial use.
K.2 Prompt
We generate the summary of climate report using following prompt on GPT4.
You are a professional weather forecast message writer. You are provided with the weather
forecasting results in the next 6 hours and you should transcribe it as readable text.The
weather forecasting results are given in json string format. One is the coarse grained
weather of the next 6 hours and the fine grained json string contains the weather forecast
every half hour in these 6 hours.
You are suppose to summarise the weather forecast message in the following aspects, each
aspect should be a sentence or a phrase:
1. Time of Day, Month, e.g. "It’s the early morning of a day in December."
2. Current overall Weather Condition, you may use the term in coarse grained information,
e.g. "The current weather is clear."
3. Weather Trend in the 6 hours, you may summarise this according to the fine grained
information, e.g. "The weather is expected to remain clear." / "Rain is expected soon."
4. Temperature Trend in the 6 hours, you may summarise this according to the fine grained
information, e.g. "The temperature is showing a mild drop."
5. Wind Speed and Direction, you may summarise this according to the fine grained
information, e.g. "There is Light Breeze from NNW."
6. Atmospheric Pressure, describe the pressure of the atmosphere, e.g. "The atmospheric
shows very Low Pressure."
7. Humidity, describe the humidity of the atmosphere, e.g. "The humidity is very high."
The summary do not have to be very detailed, but should be clear and concise.
Note that, during summarization, you should NOT include the exact values of the weather
forecast, but only the trends and conditions. You should follow the following ranking
instructions:
1. Time of Day: $0$ 06:00 -> Early Morning, 06:00 - 12:00 -> Morning, 12:00 - 18:00
-> Afternoon, 18:00 - 24:00 -> Evening
2. Wind Direction: you should convert the wind direction of degrees to N, E, S, W, NE, SE,
SW, NW, NNE, ENE, SSE, WSW, NNW, ESE, SSW, WNW.
3. Wind Speed: you should convert the wind speed to: Less than 20 km/h -> Light Breeze,
20 to 29 km/h -> Gentle Breeze, 30 to 39 km/h -> Moderate Breeze, 40 to 50 km/h -> Fresh
Breeze, 51 to 62 km/h -> Strong Breeze, 63 to 74 km/h -> High Wind, 75 to 88 km/h -> Gale,
89 to 102 km/h -> Strong Gale, Over 102 km/h -> Storm.
4. Atmospheric Pressure: you should convert the atmospheric pressure to: (<990 mbar)
-> Very Low Pressure, (990-1009 mbar) -> Low Pressure, (1010-1016 mbar) -> Average
Pressure, (1017-1030 mbar) -> High Pressure, (>1030 mbar) -> Very High Pressure.
5. Humidity: you should convert the humidity to: (<30%) -> Very Dry, (30-50%) -> Dry,
(51-70%) -> Average Humidity, (71-90%) -> Humid, (>90%) -> Very High Humid. You
may change this to more oral expression.
6. Trend: you may use "increase", "decrease", "remain", "steady", "Go up/down"...
to describe the trend of the weather condition, temperature, wind speed,
atmospheric pressure, and humidity. You may change the expression.
Note that, the unit of the weather forecast may not provided, you should use
the following units:
 Temperature: Celsius, Wind Speed: km/h, Atmospheric Pressure/barometer: mbar,
 Humidity: %, wind direction: degree from 0 to 360 with 0 as North.
 Following are some examples of the input and output of the task, you can make
 slightly changes to the output to make it more natural and fluent, but keep the
 main information, concise and the ranking instructions in mind:
 Example Input 1:
 xxxxxxx
 Example Output 1-1:
 It’s the early morning of a day in January.
 The current weather is clear.
 The weather is expected to remain clear.
 The temperature is showing a mild drop.
 There is Light Breeze from NNW.
 The atmospheric shows Average Pressure.
 The humidity is very high.
 Example Output 1-2:
 It’s the early morning of a day in January.
 The current weather is clear.
 The weather will keep clear.
 The temperature is dropping mildly.
 There is Light Breeze from NNW.
 The atmospheric pressure is average.
 The air is very humid.
K.3 Channel Details
The meaning of each channel are as follows. We ask GPT4 to generate this document. There might be incorrectness. These sentences are then embedded as channel description embeddings used in the dataset.
@@ 5. Humidity: you should convert the humidity to: (<30%) -> Very Dry, (30-50%) -> Dry,
K.4 Dataset Structure
Weather-Captioned/
|-- caption_emb/   // Pre-embeddings
|   |-- somehash0.npy
|   |-- somehash1.npy
|-- caption_emb_large/
|   |-- somehash0.npy
|   |-- somehash1.npy
|   |-- somehash2.npy
|   |-- somehash3.npy
|-- utilities/    // utilities for spider, cleaning, embedding
|   |-- spider.ipynb
|   |-- data_clean.ipynb
|-- raw_data/   // raw weather report data
|   |-- daily_weather20xx.csv
|   |-- 6hourly_weather20xx.csv
|-- captioned_data/   // GPT captioned sentences
|   |-- weather_messages20xx.csv
|-- hashtable.json      // hash table for sentences and its hash
|-- hashtable_large.json
|-- weather_2014-18.parquet     // time series data
|-- weather_large.parquet
Appendix L Steam Details
L.1 Data source
We are not directly publishing this dataset because of the intellectual property restrictions.
For those who are interested in reproducing the dataset: We directly crawl all the event data on Steam News of each Game. We may release the script for this later.
As for the online players, we use a randomly picked publicly available database that provide the downloading of csv file containing online gamer number. We will not release any data or tools to get these data.
Appendix M Implementation Details and Hyper-Parameters
@@
Except the performance on the Weather-Captioned Dataset, all other experiments are ran on the MiniLM Embedding.
Further detailed hyperparameter settings are provided in the training scripts in our codebase. We did not perform comprehensive hyper-parameter tuning because of the constraint of compute power. Thus, we may report a sub-optimal result of TGForecaster.

### v3

@@ Appendix H Performance Visualization on Atmospheric Physics Dataset
In creating a IATSF dataset, it is advisable to avoid directly generating the description out of the forecasting horizon time series pattern as news messages, as this could lead to information leakage. News messages should instead contain relevant, known information from other sources. Thus, we get the weather time series data from: https://www.bgc-jena.mpg.de/wetter/ and weather report from https://www.timeanddate.com/weather/germany/jena/historic. We will publish this dataset with CC BY-NC-SA 4.0 license since the data source forbids commercial use.
P.4.2 Motivation
The Atmospheric Physics dataset is designed as a semi-controlled environment to rigorously test the model’s ability to learn causal relationships between text and time series, as well as its text-guided channel independence and generalizability.
Such scenarios are commonly encountered in industrial applications, where correlated text and time series data often coexist. However, obtaining and releasing industrial datasets is challenging due to intellectual property restrictions. To address this, we chose the weather system—a widely available, well-understood, and publicly accessible domain—to simulate these scenarios.
As an off-the-shelf IATSF dataset, the Atmospheric Physics dataset provides a benchmark for evaluating the model’s capacity to learn causal relationships between text and time series patterns, offering a practical and accessible alternative for research and experimentation.
P.4.3 Statistical detail of the Time Series
For better understanding of the statistical distribution of Weather dataset, we plot the histogram of all 21 channels in Fig. 13.
Figure 13: Histogram of all 21 channels. It shows all the channels have unique value distribution, making a model hard to generalize on all channels without knowing related information.
Table 13: Example caption of the Atmospheric Physics dataset.
Topic Example
Month & Time of the Day It’s the early morning of a day in January.
Overall Weather The current weather is clear.
Weather Trend in next 6h The weather is expected to remain clear.
Temperature Trend in next 6h The temperature is showing a mild drop.
Wind Speed & Direction There is Light Breeze from NNW.
Atmosphere Pressure Level The atmospheric shows Average Pressure.
Humidity Level The air is very humid.
P.4.4 Use of Large Language Models for Preprocessing the Weather Dataset
We would like to clarify that the use of Large Language Models (LLMs) in this work is strictly limited to the preprocessing and creation of the Atmospheric Physics dataset. LLMs are not part of our model or method, nor do they contribute to the training or inference process of FIATS. The Atmospheric Physics dataset is intended to serve as an off-the-shelf, text-time synchronized benchmark dataset with raw text and pre-embedded text embeddings as optional inputs.
The primary reason for using LLMs to preprocess this dataset is to generate diverse and correlated textual descriptions, ensuring a richer corpus for training and evaluation. By incorporating varied expressions, we enable the model to generalize to different textual forms while aligning the semantic meaning of text with time series patterns. For instance, the descriptions “The morning will be sunny, but clouds will increase in the afternoon with a chance of light rain” and “The day starts with clear skies, gradually turning cloudy with some rain in the afternoon” carry the same semantic information but differ in expression. This diversity enhances the robustness of the benchmark and validates the model’s generalization capabilities.
Additionally, the raw data source for this dataset often includes general weather reports in text form, accompanied by coarse numerical updates every six hours. While numerical values such as {High_Temp: 25, Low_Temp: 20, Temp_Trend: slightly increasing, Wind_Speed: 5, Wind_Direction: East} are available, they lack the precision required for reliable exogenous variables. Moreover, the raw text contains rich semantic details—such as qualitative weather descriptions—that cannot be effectively captured using numerical values or one-hot encoding. Using text embeddings allows the model to leverage both semantic and numerical information more effectively.
In summary, the LLM preprocessing step is solely for dataset preparation and corpus diversity, ensuring that the Atmospheric Physics dataset is suitable for evaluating text-guided time series forecasting models. Our method does not rely on any LLM capabilities, and the inclusion of LLM-generated text is not a necessary step for IATSF or any similar model. We will include raw data samples in the final paper to provide greater clarity and avoid any misunderstandings.
P.4.5 Channel Details
The meaning of each channel are as follows. The original weather dataset only contains the abbreviation for each channel, to further enrich the semantic for accurate information, we add a line of explanation after it as the channel description.
@@
P.4.6 Visualization of the Test Sample
We show a segment of test sample along with the dynamic news timeline in Fig. 14. The news messages are sparse and vague and not directly correlated to some of the channels. These text are passed to the model as text embeddings and aligned with time series on time domain. Thus, the model can extract causal relationship to guide each channel to perform accurate prediction even though they have distinguished distribution.
Figure 14: An visualization of test sample with all the corresponding dynamic news. Atmospheric Physics dataset have dynamic weather report update every 6 hours. As we demonstrate a case of predicting 48 hours. Note that the embedding of these sentences are fed to the FIATS along with the look-back window time series as input. We highlight some of the words that may make impact on the forecasting result.
The following sections give detailed performance and visualization across all the channels.
P.5 GAUD Details
GAUD datasets contains 89 subdataset. Each subdataset correspond to the active user time series of one specific game along with text information. Each subdataset contains one basic information as the channel description includes game title, genera and developer also with the update log includes release date, update type, update title, and article body. We will release the pre embeddings of these text information to avoid violating intellectual property constrains.
Appendix Q Implementation Details and Hyper-Parameters
@@
Except the performance on the Atmospheric Physics Dataset, all other experiments are ran on the MiniLM Embedding. We selected MiniLM as the embedding model because it achieves results comparable to OpenAI embeddings while producing smaller embeddings (384 dimensions for MiniLM versus 512 for OpenAI). This reduced embedding size speeds up training, particularly for ablation studies, making it more practical for our experiments.
Further detailed hyperparameter settings are provided in the training scripts in our codebase. We did not perform comprehensive hyper-parameter tuning because of the constraint of compute power. Thus, we may report a sub-optimal result of FIATS.